find the exact value of sin2ø, cos2ø,tan2ø, and the quadrant in which 2ø lies.
cosø=-33/65, ø in quadrant 3rd

Answer by Alan3354(69443)

find the exact value of sin2ø, cos2ø,tan2ø, and the quadrant in which 2ø lies.
cosø=-33/65, ø in quadrant 3rd
----
sinø = sqrt(1 - cos^2) = -56/65 *** negative in Q3
---
sin2ø = 2sinø*cosø = 56*33/65^2 = 1848/4225
---
cos2ø = sqrt(1-sin^2) = -sqrt((4225^2 - 1848^2)/4225^2) = -sqrt(14435521)/4225
---
tan2ø = sin/cos = -1848/sqrt(14435521)= -1848*sqrt(14435521)/14435521
---
2ø is in Q2